OpenStudy (anonymous):

Plot this complex number in the complex plane and write it in polar form. Express the argument in degrees. 5 + 2i Give your answer using the form z = A(cos(B°) + i sin(C°)), where A > 0; 0° ≤ B° < 360°; and 0° ≤ C° < 360°. Round to 2 decimal places. A = B = C =

OpenStudy (amistre64):

5 over and 2 up; tan^-1(2/5) is the angle

OpenStudy (amistre64):

r = the absolute value of the complex number; the distance from the origin sqrt(25+4) = sqrt(29) = r

OpenStudy (amistre64):

a = sqrt(29) = 5.385...

OpenStudy (amistre64):

b = 21.80.... and c = 21.80... i believe

OpenStudy (amistre64):

why they give you b and c as different i dunno, but thhey are the same angle

OpenStudy (anonymous):

B= C= ? who writes this?